# From Spurious to Causal: Low-rank Orthogonal Subspace Intervention for Generalizable Face Forgery Detection

### v1

@@ 4.3 How to remove spurious correlation subspaces and why our method works
We define spurious correlation as representation directions that correlate with the label via forgery-irrelevant information (e.g., background, identity) rather than forgery traces. Formally, letting the representation be $R=h(Z_{s},Z_{c},\varepsilon)$ and taking a first-order local expansion around a reference point yields $R\approx J_{s}\,\Delta z_{s}+J_{c}\,\Delta z_{c}$, where the variation aligned with $Z_{s}$ (the spurious component $N:=J_{s}\,\Delta z_{s}$) is considered to be low-rank and approximately additive with the forgery-related component $S:=J_{c}\,\Delta z_{c}$. We model the spurious correlation subspace by a column-orthonormal basis $Q\in\mathbb{R}^{D\times r}$ with $r\ll D$, obtained by orthogonalizing a slim learnable matrix $B$ via QR decomposition to get $Q$, which induces the projector $P=QQ^{\top}$ and its orthogonal complement $P_{\perp}=I-P$. At both training and inference, we remove spurious correlations by applying $P_{\perp}$ to token features. Under the row-wise convention for features $X\in\mathbb{R}^{N\times D}$, we compute $X_{c}=XP_{\perp}=X(I-P)$, thus preserving forgery-related signal while suppressing nuisance directions. The orthogonal basis $Q$ is learned end-to-end purely from the cross-entropy loss: if $Q$ captures shortcuts, projecting them out improves generalization; if it mistakenly captures causal factors, the loss penalizes it and drives $Q$ away. This yields a precise mechanism to excise spurious correlation subspaces without any explicit supervision of $Z_{s}$ or $Z_{c}$.
@@
[…]
Table 1: Cross-dataset evaluation results (Frame-level AUC). All methods are trained on FF++ and evaluated on other datasets. The best results are indicated in bold and the second-best results are underlined. “” indicates that we reproduce results with its original code.
@@
[…]
Table 2: Cross-dataset evaluation results (Video-level AUC). All the results are taken from their original papers.
@@
[…]
Table 3: Cross-dataset evaluation results (Video-level AUC). Comparing our method with other CLIP-based methods.
@@
[…]
Table 4: Cross-manipulation evaluations within FF++ domain. All models are trained on FF++ (c23) and tested on the latest DF40[46] dataset. The best results are indicated in bold and the second-best results are underlined.
5 Experiments
5.1 Experimental Settings
Datasets. To comprehensively and fairly evaluate the model’s generalization ability, we follow the DeepFakeBench protocol [48] and utilize six publicly available datasets: FaceForensics++ (FF++)[40], Celeb-DF (CDF-v1 and CDF-v2)[23], Deepfake Detection Challenge dataset (DFDC)[9], DeepFake Detection Challenge Preview (DFDCP)[8], and Deepfakedetection (DFD)[7]. Our method is trained on c23 compression version of FF++ and tested on other datasets. In this paper, following the latest method [45, 49], we adopt two widely used and standard protocols for evaluation: Protocol-1: cross-dataset evaluation, Protocol-2: cross-manipulation evaluation within the FF++ data domain. For Protocol-1, all methods are trained on FF++ c23 training set and evaluated on other deepfake datasets. For Protocol-2, we evaluate the models on the latest deepfake dataset DF40 [46], which includes four types of forgery methods: face swapping, face reenactment, whole face synthesis, and face editing. Moreover, these forged faces are generated within the FF++ domain, ensuring the manipulation methods different while the data domains remain unchanged. For specific experimental details, refer to the supplementary material Sec.4.
@@ 5.1 Experimental Settings
Table 5: Ablation study on visual space intervention. Evaluation metrics are the frame-level AUC, AP, and EER, respectively.
@@
Table 6: Ablation study on rank of spurious correlation subspace and number of intervention layers. Evaluation metrics are the frame-level AUC, AP, and EER, respectively.
Figure 3: Robustness Analysis. Our method is compared with CLIP[38], SRM[29], and CORE[34] across five levels of six particular types of perturbations in video-level AUC.
5.2 Comparison with State-of-the-art Methods
In this section, we compare our approach with other methods under both frame-level and video-level evaluations.
Frame-level Evaluation. Table 1 shows the cross-dataset evaluation results in frame-level AUC. The results for ED [1], CFM [28], and FIA-USA [30] are directly cited from the original paper and other results are cited from Forensics-Adapter [6] under DeepFakeBench [48]. Effort [45] do not report frame-level results, so we reproduce them using its open source code. From the table, we can observe that our approach outperforms the current state-of-the-art methods Effort and Forensics-Adapter. On the widely regarded most challenging DFDC and DFDCP datasets, our method surpasses Forensics-Adapter by 1% (from 0.843 to 0.853) and 1.5% (from 0.890 to 0.905) AUC respectively. These results demonstrate the excellent generalization performance of our method.
Video-level Evaluation. Table 2 shows the cross-dataset evaluation results in video-level AUC. We compare 10 state-of-the-art methods, with all results directly cited from their original papers. In video-level comparisons, our approach outperform the state-of-the-art Effort by 3.4% (from 0.843 to 0.877) and 2% (from 0.909 to 0.929) AUC on the DFDC and DFDCP datasets respectively, providing compelling evidence of its superior generalization.
Compared with CLIP-based Methods. Table 3 presents a comparison of our method with other CLIP-based methods. It can be observed that our method outperforms the second-best approach by 1.6% (from 0.861 to 0.877) and 2% (from 0.909 to 0.929) on the DFDC and DFDCP datasets respectively. It is noteworthy that our approach significantly outperforms UDD, which specifically distinguishes and addresses one particular type of spurious correlation. The results further demonstrate the effectiveness of our approach.
Cross-manipulation Evaluation. Table 4 shows cross-manipulation evaluation results within FF++ domain. All models in the table were trained on the FF++ dataset and then tested on the DF40 [46] dataset. Note that DF40 includes multiple manipulation techniques (e.g., face swapping, face reenactment). It can be observed that our method achieves excellent detection performance when confronted with different manipulation techniques. In particular, our method outperforms the second-best method by 6.5% (from 0.875 to 0.940) and 3.6% (from 0.844 to 0.880) on the e4s [27] and FaceDancer [39] datasets, respectively.
5.3 Ablation Study
Effectiveness of LROR. To verify that our method can project spurious correlations into a low-rank subspace and remove them from the original representation, we conduct a counterfactual validation experiment. Quantitative and qualitative experiments are shown in Table 5 and Figure 4. We replace the forward visual space with two complementary forms for comparative validation. The first is the non-causal factor space $\mathbf{X_{vis}}\mathbf{Q}\mathbf{Q}^{\top}$ (denoted by SP, representing the subspace of spurious correlations). We achieve this setting by fixing $\mathbf{Q}$ obtained from trained SeLop, replacing only the forward flow with $\mathbf{X_{vis}}\mathbf{Q}\mathbf{Q}^{\top}$, and retraining the linear classification head. The second is the causal factor space $\mathbf{X_{vis}}(\mathbf{I}-\mathbf{Q}\mathbf{Q}^{\top})$ (denoted by CA, representing the visual space after our method’s intervention). ID1 in Table 5 is the Vanilla CLIP, without subspace replacement or intervention, where forgery-related knowledge is intertwined with irrelevant factors to forgery. As can be seen in Figure 4 (a), the distributions of real and forged samples overlap significantly, making it difficult to form a robust decision boundary. When the visual space consists solely of non-causal factors (corresponding to ID2 and Figure 4 (b)), the AUC performance across several datasets hovers around 0.5, with lower AP and higher error rate. This near-random performance suggests that this subspace primarily contains non-causal features, making it difficult for the model to accurately distinguish real from fake faces based on this information. However, after intervention with our method (corresponding to ID3 and Figure 4 (c)), the AUC and AP on each benchmark significantly improve, the EER significantly decreases, and T-SNE visualizations [31] demonstrate clear between-class separation. This demonstrates that our method effectively removes the low-rank spurious correlation subspace induced by confounding factors at the representation level and focuses the spurious correlation factors on orthogonal complements, resulting in consistent improvements across datasets.
Impact of Different Hyperparameters. In this section, we investigate the impact of different ranks of the low-rank subspace and numbers of intervention layers on detection performance of our model. As shown in the Table 6, we can observe that: (1) For the rank of the spurious correlation space, if rank is too small, it fails to eliminate most spurious correlation factors, and if rank is too large, it may hinder the sufficient learning of forgery-related features. (2) Regarding the number of intervention layers, fewer intervention cannot completely decouple spurious correlation in the visual space and excessive intervention may destroy pre-trained knowledge of the CLIP [38]. Therefore, we select rank=32 and layers=12 as the hyperparameters for the model in this paper.
@@ 5.3 Ablation Study
Complexity Analysis. In this section, we examine the computational complexity of our approach. By freezing all parameters of CLIP and only training the low-rank orthogonal space for visual intervention alongside a linear classification head, our model requires a mere 0.43 million trainable parameters. Compared to methods employing adapters for fine-tuning, our approach is remarkably lightweight. As demonstrated in Figure 5, our approach achieves the best performance with the fewest trainable parameters, thereby proving its superiority.
6 Conclusion
In this paper, we analyze the reasons of failures of Vanilla CLIP’s generalization from the perspective of causal learning. Unlike previous approaches that identify specific spurious correlation and then design specialised modules to mitigate their effects, we treat spurious correlations as a collective phenomenon. Specifically, we intervene in CLIP’s visual representation space by removing spurious correlation subspaces through orthogonal low-rank projection, compelling the model to implement classification based on real causal factors (e.g., forgery-related features). Qualitative and quantitative analysis demonstrated its effectiveness. Extensive experiments on several benchmarks demonstrate our approach’s outstanding generalization and robustness.

### v2

@@ III-C Why our method works
At both training and inference, we remove spurious correlations by applying $P_{\perp}$ to token features. Under the row-wise convention for features $X\in\mathbb{R}^{N\times D}$, we compute $Z_{c}=XP_{\perp}=X(I-P)$, thus preserving forgery-related signal while suppressing spurious directions. The orthogonal basis $Q$ is learned end-to-end purely from the cross-entropy loss: if $Q$ captures shortcuts $Z_{s}$, projecting them out improves generalization; if it mistakenly captures causal factors $Z_{c}$, the loss penalizes it and drives $Q$ away. This yields a precise mechanism to excise spurious correlation subspaces without any explicit supervision of $Z_{s}$ or $Z_{c}$.
IV Experiments
In this section, we first introduce the datasets used in the subsequent experiments. Next, we describe the implementation details of SeLop. We then discuss the experimental results of generalization evaluation, including cross-dataset evaluation, cross-manipulation method evaluation to demonstrate the effectiveness of the proposed method. Moreover, we discuss the robustness of SeLop under different perturbations. Additionally, we conduct ablation experiments to validate the effectiveness of LROR by qualitative and quantitative analysis. Finally, we conduact hyperparameter sensitivity analysis and complexity analysis of SeLop.
IV-A Experimental Settings
Datasets. To comprehensively and fairly evaluate the model’s generalization ability, we follow the DeepFakeBench protocol [64] and utilize six publicly available datasets: FaceForensics++ (FF++)[48], Celeb-DF (CDF-v1 and CDF-v2)[28], Deepfake Detection Challenge dataset (DFDC)[11], DeepFake Detection Challenge Preview (DFDCP)[10], and Deepfakedetection (DFD)[7].
IV-A1 FaceForensics++
FaceForensics++ [48] is a standard benchmark in face forgery detection. It includes 1,000 real videos and 4,000 fake videos, which consist of four types of face manipulation techniques: Deepfakes (DF) [8], Face2Face (F2F) [54], FaceSwap (FS) [24], and NeuralTextures (NT) [53]. In our experiments, we use the c23 compression version of FF++ to train and evaluate all the methods.
IV-A2 Celeb-DF
Celeb-DF-v1 [28] includes 408 real videos and 795 synthesized videos with reduced visual artifacts. Celeb-DF-v2 [28] is a larger dataset forged by several forgery techniques and is an extension of the Celeb-DF-v1 version. It consists of 590 real videos, featuring subjects of different ages, races, and along with 5639 corresponding DeepFake videos.
IV-A3 Deepfake Detection Challenge dataset
DFDC [11] is a challenging large-scale video dataset, since its real videos are very close to real life, and the forgery traces are smaller and more subtle. It contains over 100,000 videos, so we only conduct cross-dataset evaluation on its test set. DFDCP [10] is a dataset of 5,000 videos created by Facebook to advance deep forgery detection technology, which is the preview version of the Deepfake Detection Challenge (DFDC) dataset.
IV-A4 DeepFakeDetection
DFD [7] consists of 363 real videos and 3071 synthetic videos, which is generated by Google and collected in FF++ now.
In this paper, we adopt four widely used and standard protocols for evaluation: Protocol-1: cross-dataset evaluation, Protocol-2: cross-manipulation evaluation on the latest DF40 and on FF++ (c23). Protocol-3: Intra-dataset evaluation. Protocol-4: detection in real-world scenarios. For Protocol-1, all methods are trained on FF++ c23 training set and evaluated on other deepfake datasets. For Protocol-2, on one hand, we evaluate the models on the latest deepfake dataset DF40 [62], which includes four types of forgery methods: face swapping, face reenactment, whole face synthesis, and face editing. On the other hand, we train models on one manipulation type of FF++ and test them on three unknown forgery techniques. For Protocol-3, all methods are trained on Celeb-DF-v2 dataset and evaluated on it too. For Protocol-4, we evaluate our model on latest datasets, DDL [40]. It offers more diverse forgery methods and facial scenes, serving as more challenging benchmarks for real-world scenarios.
TABLE I: Cross-dataset evaluation results (Frame-level AUC). The best results are indicated in bold and the second-best results are underlined. “” indicates that we reproduce results with its original code.
@@ IV-A4 DeepFakeDetection
SeLop (Ours) - 0.924 0.901 0.853 0.905 0.926 0.902
TABLE II: Cross-dataset evaluation results (Video-level AUC). All the results are taken from their original papers.
@@
SeLop (Ours) - 0.948 0.877 0.929
TABLE III: Cross-dataset evaluation results (Video-level AUC). Comparing our method with other CLIP-based methods.
@@
SeLop (Ours) - 0.948 0.877 0.929
Implementation Details. Our method is implemented by Pytorch 1.12.0 with one Nvidia GTX 4090 GPU. We utilize CLIP ViT-L/14 [46] as the default vision foundation model. In subsequent experiments, we also verify the effect on other versions of CLIP. Following the configuration of DeepFakeBench [64]: For each video, 32 frames were extracted for training or testing, and RetinaFace [9] was used for face extraction. We use the fixed learning rate of 2e-4 and weight decay of 0.0005 for training our model on FF++ [48] c23 training set. Adam optimiser is employed for optimization. We set the batch size to 32 for both training and testing. In addition, the rank for the low-rank spurious correlation subspace is set to 32 and the last 12 layers serve as the target layers for our intervention. To ensure comprehensive and fair comparisons, all our experiments were conducted under the DeepFakeBench framework [64].
IV-B Comparison with State-of-the-art Methods
In this section, we conduct cross-dataset and intra-dataset experiments to evaluation our method and compare SeLop with other SOTA methods.
Protocol-1: Cross-dataset Evaluation. We report frame-level and video-level results. Furthermore, we compare our method with state-of-the-art CLIP fine-tuning based methods to demonstrate the superiority of SeLop low-rank interventions.
IV-B1 Frame-level Evaluation
Table I shows the cross-dataset evaluation results in frame-level AUC. The results for ED [1], CFM [34], and FIA-USA [36] are directly cited from the original paper and other results are cited from Forensics-Adapter [6] under DeepFakeBench [64]. Effort [61] do not report frame-level results, so we reproduce them using its open source code. From the table, we can observe that our approach outperforms the current state-of-the-art methods Effort and Forensics-Adapter. On the widely regarded most challenging DFDC and DFDCP datasets, our method surpasses Forensics-Adapter by 1% (from 0.843 to 0.853) and 1.5% (from 0.890 to 0.905) AUC respectively. These results demonstrate the excellent generalization performance of our method.
TABLE IV: Cross-manipulation evaluations (Video-level AUC). We evaluate all methods on the latest DF40[62] dataset. The best results are indicated in bold and the second-best results are underlined.
@@ IV-B1 Frame-level Evaluation
MAP-Mamba [56] 0.815 0.896 0.959 0.779 0.826 - 0.855 0.813 -
SeLop (Ours) 0.925 0.866 0.957 0.940 0.880 0.952 0.897 0.856 0.909
IV-B2 Video-level Evaluation
Table II shows the cross-dataset evaluation results in video-level AUC. We compare 10 state-of-the-art methods, with all results directly cited from their original papers. In video-level comparisons, our approach outperform the state-of-the-art Effort by 3.4% (from 0.843 to 0.877) and 2% (from 0.909 to 0.929) AUC on the DFDC and DFDCP datasets respectively, providing compelling evidence of its superior generalization.
IV-B3 Compared with CLIP-based Methods
Table III presents a comparison of our method with other CLIP-based methods. It can be observed that our method outperforms the second-best approach by 1.6% (from 0.861 to 0.877) and 2% (from 0.909 to 0.929) on the DFDC and DFDCP datasets respectively. It is noteworthy that our approach significantly outperforms UDD, which specifically distinguishes and addresses one particular type of spurious correlation. The results further demonstrate the effectiveness of our approach.
TABLE V: Cross-manipulation evaluation results (Video-level AUC). Cross Avg. represents the average results on three cross-manipulation evaluation trials.
Methods Train DF F2F FS NT Cross Avg.
EN-B4 [52] DF 0.999 0.683 0.452 0.657 0.597
RECCE [3] 0.999 0.698 0.547 0.772 0.672
CFM [34] 0.999 0.776 0.549 0.750 0.692
SeLop 0.999 0.753 0.980 0.637 0.790
EN-B4 [52] F2F 0.818 0.991 0.583 0.665 0.688
RECCE [3] 0.716 0.992 0.500 0.723 0.646
CFM [34] 0.819 0.992 0.601 0.708 0.709
SeLop 0.948 0.997 0.907 0.630 0.828
EN-B4 [52] FS 0.682 0.669 0.996 0.512 0.621
RECCE [3] 0.631 0.662 0.997 0.581 0.624
CFM [34] 0.729 0.714 0.999 0.517 0.653
SeLop 0.993 0.839 0.999 0.565 0.799
EN-B4 [52] NT 0.821 0.750 0.493 0.991 0.688
RECCE [3] 0.724 0.647 0.516 0.996 0.629
CFM [34] 0.883 0.768 0.526 0.992 0.726
SeLop 0.895 0.820 0.711 0.979 0.809
TABLE VI: Intra-dataset comparison in terms of ACC and AUC. Comparison with State-of-the-art methods on Celeb-DF-v2 dataset. The best results are indicated in bold and the second-best results are underlined.
Method CelebDF-v2
ACC AUC
F3-Net [45] 0.960 0.989
Two-branch [38] 0.938 0.959
TD3DCNN [69] 0.811 0.888
SRM [35] 0.883 0.967
PEL [15] 0.968 0.982
STDT [70] 0.917 0.972
CORE [42] 0.854 0.992
PCC [19] - 0.990
LPS [14] - 0.968
SPSL [32] 0.916 0.996
MADD [72] 0.946 0.980
SeLop 0.968 0.999
TABLE VII: Face Forgery Detection in real-world scenes on the latest DDL [40] dataset. We compare SeLop with eight SOTA methods.
Method DDL
ACC AUC
Xception [5] 0.435 0.538
F3-Net [45] 0.453 0.523
RFM [55] 0.523 0.573
SRM [35] 0.560 0.503
CORE [42] 0.443 0.535
SPSL [32] 0.607 0.613
MAP-Mamba [56] 0.672 0.712
ForensicAdapter [6] 0.779 0.835
SeLop 0.855 0.933
Figure 4: Robustness Analysis. Our method is compared with CLIP[46], SRM[35], CORE[42], ForensicAdapter [6], Effort[61] across five levels of six particular types of perturbations in video-level AUC.
Protocol-2: Cross-manipulation Evaluation
Table IV shows cross-manipulation evaluation results. All models in the table were trained on the FF++ dataset and then tested on the latest DF40 [62] dataset. Note that DF40 includes multiple manipulation techniques (e.g., face swapping, face reenactment). It can be observed that our method achieves excellent detection performance when confronted with different manipulation techniques. In particular, our method outperforms the second-best method by 6.5% (from 0.875 to 0.940) and 3.6% (from 0.844 to 0.880) on the e4s [33] and FaceDancer [47] datasets, respectively.
In addition, we conduct cross-manipulation experiments on FF++ (C23) to evaluate the detector’s generalization to unknown manipulation methods. We train the model on one forgery type and evaluate it on all four types (DF, F2F, FS, and NT). As shown in Table V, when the models are trained on the FS type, SeLop outperforms CFM in cross-manipulation evaluations with a near 15% Cross Avg. AUC improvement. This strongly demonstrates that our method does not overfit to any particular forgery pattern, but instead learns a general representation.
Protocol-3: Intra-dataset Evaluation
To further demonstrate the effectiveness of our method on the intra-dataset, we compare it with other state-of-the-art methods on Celeb-DF-v2 [28]. As shown in Table VI, we can see that SeLop outperforms all methods and achieves an impressive AUC performance of 0.999. This demonstrates SeLop’s powerful ability to learn what is “fake”.
Protocol-4: Detection in real-world scenarios
It is an important protocol in face forgery detection to evaluate models on benchmarks oriented towards real-world scenarios. As shown in Table VII, it can be observed that our method outperforms all methods in terms of ACC and AUC. Surprisingly, SeLop achieves an image-level AUC of 0.933 on DDL dataset [40].
TABLE VIII: Ablation study on visual space intervention. Evaluation metrics are the frame-level AUC, AP, and EER, respectively.
@@ IV-B3 Compared with CLIP-based Methods
IV-C Ablation Study
Effectiveness of LROR. To verify that our method can project spurious correlations into a low-rank subspace and remove them from the original representation, we conduct a counterfactual validation experiment. Quantitative and qualitative experiments are shown in Table VIII and Figure 5. We replace the forward visual space with two complementary forms for comparative validation. The first is the non-causal factor space $\mathbf{X_{vis}}\mathbf{Q}\mathbf{Q}^{\top}$ (denoted by SP, representing the subspace of spurious correlations). We achieve this setting by fixing $\mathbf{Q}$ obtained from trained SeLop, replacing only the forward flow with $\mathbf{X_{vis}}\mathbf{Q}\mathbf{Q}^{\top}$, and retraining the linear classification head. The second is the causal factor space $\mathbf{X_{vis}}(\mathbf{I}-\mathbf{Q}\mathbf{Q}^{\top})$ (denoted by CA, representing the visual space after our method’s intervention). ID1 in Table VIII is the Vanilla CLIP, without subspace replacement or intervention, where forgery-related knowledge is intertwined with irrelevant factors to forgery. As can be seen in Figure 5 (a), the distributions of real and forged samples overlap significantly, making it difficult to form a robust decision boundary. When the visual space consists solely of non-causal factors (corresponding to ID2 and Figure 5 (b)), the AUC performance across several datasets hovers around 0.5, with lower AP and higher error rate. This near-random performance suggests that this subspace primarily contains non-causal features, making it difficult for the model to accurately distinguish real from fake faces based on this information. However, after intervention with our method (corresponding to ID3 and Figure 5 (c)), the AUC and AP on each benchmark significantly improve, the EER significantly decreases, and T-SNE visualizations [37] demonstrate clear between-class separation. This demonstrates that our method effectively removes the low-rank spurious correlation subspace induced by confounding factors at the representation level and focuses the spurious correlation factors on orthogonal complements, resulting in consistent improvements across datasets.
TABLE IX: Ablation study on rank of spurious correlation subspace and number of intervention layers. Evaluation metrics are the frame-level AUC, AP, and EER, respectively.
@@ IV-C Ablation Study
TABLE X: Influence of intervention on various CLIP architectures. (Frame-level AUC).
Method CDF-v2 DFDC DFD
CLIP ViT-B/32 [46] 0.662 0.662 0.764
[…]
CLIP ViT-B/16 [46] 0.721 0.693 0.795
[…]
CLIP ViT-L/14 [46] 0.763 0.742 0.828
[…]
Impact of Different Hyperparameters. In this section, we investigate the impact of different ranks of the low-rank subspace and numbers of intervention layers on detection performance of our model. As shown in the Table IX, we can observe that: (1) For the rank of the spurious correlation space, if rank is too small, it fails to eliminate most spurious correlation factors, and if rank is too large, it may hinder the sufficient learning of forgery-related features. (2) Regarding the number of intervention layers, fewer intervention cannot completely decouple spurious correlation in the visual space and excessive intervention may destroy pre-trained knowledge of the CLIP [46]. Therefore, we select rank=32 and layers=12 as the hyperparameters for the model in this paper.
Figure 6: The visualization of GradCAM [50] on the Vanilla CLIP [46] and our method (cross-manipulation). The eight images in the top row correspond to fake images generated by the four forgery techniques, including DeepFake [8], Face2Face [54], FaceSwap [24], and NeuralTextures [53]. The following two rows show the attention maps for Vanilla CLIP and our method respectively.
Figure 7: The visualization of GradCAM [50] on the Vanilla CLIP [46] and our method (cross-datasets). The eight images in the top row correspond to fake images from four standard benchmarks, including Celeb-DF-v2 [28], DFD [7], DFDC [11], and DFDCP [10]. The following two rows show the attention maps for Vanilla CLIP and our method respectively.
@@
Effects of intervention on different CLIP architectures. In this section, we investigate the impact of visual-spatial interventions on different CLIP [46] architectures, including CLIP ViT-B/32, CLIP ViT-B/16, and CLIP ViT-L/14. Our approach utilize CLIP ViT-L/14 as backbone. As Table X clearly demonstrates, detection performance significantly improved across all CLIP variants following low-rank projection to eliminate spurious correlations. Notably, on the CDF-v2 dataset, all three architectures achieve over 12% improvement in frame-level AUC performance after intervention. This fully validates the effectiveness of our methodology.
Qualitative Results. In this section, we use Grad-CAM [50] to visualize the location of model’s attention on the faces generated by four forgery techniques (DeepFakes [8], Face2Face [54], FaceSwap [24], and NeuralTextures [53]) and four standard benchmarks (Celeb-DF-v2 [28], DFD [7], DFDC [11], and DFDCP [10]), which is shown in Figure 6, 7. It can be noticed that compared to baseline, our method better identifies and utilizes facial forgery cues to judge authenticity, greatly mitigating the negative interference of spurious correlation factors. Visualization results prove the effectiveness of our method.
IV-D Complexity Analysis
In this section, we examine the computational complexity of our approach. By freezing all parameters of CLIP and only training the low-rank orthogonal space for visual intervention alongside a linear classification head, our model requires a mere 0.39 million trainable parameters. Compared to methods employing adapters for fine-tuning, our approach is remarkably lightweight. As demonstrated in Figure 8, our approach achieves the best performance with the fewest trainable parameters, thereby proving its superiority.
V Conclusion
In this paper, we analyze the reasons for Vanilla CLIP’s generalization failure through GradCAM visualization and PCA empirical research: low-rank spurious bias. To address this, we innovatively propose a simple and effective representation space intervention paradigm from the perspective of causal representation learning. Specifically, we remove the spurious correlation subspace through orthogonal low-rank projection, forcing the model to classify based on authentic causal factors. Both qualitative and quantitative analyses validate the effectiveness of the proposed method. Extensive experiments on several benchmark datasets demonstrate that our method possesses excellent generalization ability and robustness.